5.0 balloon has a pressure of 1.0 atm at sea level, where the temperature is 28°C. The balloon was let loose and ascended to a height in the atmosphere where the temperature is 11 °C. What will be the volume of the balloon if the atmosphere pressure at the new altitude is 0. 65 atm
1
Expert's answer
2020-11-19T09:19:53-0500

We know that


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2},\\\space\\
V_2=V_1\frac{T_2P_1}{T_1P_2}=V_1\frac{(273+11)(1)}{(273+28)(0.65)}=1.45V_1."